OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1

OpenStudy (goten77):

any1 can post numbers zordoloom 2x+y=3 same thing as y=3-2x plug this into bottom equation and get -3x +4(3-2x)=1 then get -3x + 12 - 8x= 1 then -11x +12=1 then -11x=-11 then divide by -11 and get that x = 1 plug this into original and get that 2+y=3 and find out that y = 1 therefor y=1 and x=1

OpenStudy (anonymous):

A. (1, 1) B.(-1, -1) C.(1, -1) D.(-1, 1)

OpenStudy (anonymous):

Thanks.